# Newtonstr. 15, 12489 Berlin, Germany

### v1

@@ 1 Introduction
Despite these advances, several open questions remain regarding the impact of gauge degrees of freedom (DOF) truncation, the role of external fields in string dynamics, and the dynamics of the theory’s stable particle, called the Schwinger boson. Prior studies often truncate the gauge DOF Hilbert space to allow only a small number of states per link, potentially limiting the proximity to the continuum limit. Additionally, the influence of an applied background electric field on string behavior in open quantum systems (OQS) has not been explored.
@@ 2 Theoretical framework of the open lattice Schwinger model
An open quantum system is described by the system of interest $S$ and its environment $E$ through the Hamiltonians $H_{S}$, $H_{E}$ and $H_{I}$, where the latter is the interaction Hamiltonian between $S$ and $E$ Breuer:2002pc ; Rivas:2012ugu . It is usually the case that we are interested in measuring observables only on $S$ but nevertheless want to study the effect of $E$ on $S$. Hence, we only aim follow the dynamics of the reduced density matrix for the system $\rho_{S}$ defined by tracing out the degrees of freedom of $E$.
@@
where $\rho_{E}=e^{-\beta H_{E}}/\text{Tr}(e^{-\beta H_{E}})$ is the environment’s density matrix in the Gibbs state. Further, we work in the quantum Brownian motion (QBM) limit given by the separation of time scales $\tau_{R}\gg\tau_{E}$, $\tau_{S}\gg\tau_{E}$. The first inequality implies that the system with relaxation time $\tau_{R}\sim T/H_{I}^{(\text{int})^{2}}$ relaxes much more slowly than the environment’s correlation time $\tau_{E}\sim 1/T$, where $H_{I}^{(\text{int})}$ is the interaction Hamiltonian in the interaction picture. We can interpret $H_{I}^{(\text{int})}$ in the definition of $\tau_{R}$ as the gap between the ground and first excited state of $H_{I}^{(\text{int})}$. The second inequality suggests that the intrinsic time scale of the system $\tau_{S}\sim 1/H_{S}$ is much greater than $\tau_{E}$, which is valid in the limit $T\gg H_{S}$. Here we can interpret $H_{S}$ to be the spectral gap of $H_{S}$ which is $1/\sqrt{\pi}$ at $m=0$ in the continuum limit adam . These limits further justify the assumption that $\rho_{E}$ remains for all times in the thermal Gibbs state defined by $H_{E}$ as any perturbation on the environment’s Gibbs state will decay too fast to affect the system which we are interested in. The high temperature limit we are using where the environment is always in thermal equilibrium is relevant for the case of a hot nuclear QGP medium oqs_for_quarkonia . Under these assumptions, the evolution of the density matrix of $S$ is given by the Lindblad master equation
@@ 3 Lindbladian evolution scheme with tensor networks
We represent the right-hand side of Eq. (19) as a global MPO with two singular value cutoffs $\epsilon_{1}$, $\epsilon_{2}$. Using the ITensors Julia library ITensors allows to convert $\mathcal{L}_{T}$ automatically to an MPO and can apply a truncation on the singular values in the process, which we call $\epsilon_{1}$. The ratio of the sum of squares of the neglected singular values to the sum of squares of all singular values will not exceed this cutoff at any given singular value decomposition truncation. Then $\epsilon_{2}$ is used when we are multiplying $\mathcal{L}_{T}$ onto itself to form powers of this operator as needed for the right hand side of Eq. (19). Finally, we use the adaptive time-dependent DMRG algorithm Xiang_2023 to apply the above operators with a singular value cutoff $\epsilon$, and evolve our state $\rho_{S}(t)$. For more details on this algorithm and the partition of $\mathcal{L}$ into groups see appendix B. Throughout this paper we fix $D(n-k)=D\delta_{n,k}$ open_schwinger_original , $x=1$, $\epsilon=10^{-11}$, $\epsilon_{1}=\epsilon_{2}=10^{-9}$, $\kappa=2$ and $\tau=0.01$, except from section 4.4 where we set $x=4$, $\tau=0.001$ due to the larger $N$ used. The above methods allow us to simulate the time evolution of meson type states according to Eq. (14) and gather results for their dynamics in a hot medium which we turn to in the next section.
@@ 4.3 Temperature dependence of the string’s thermalization
In figure 7 we plot $\mathcal{T}$ as a function of $T$ for $D=2,5$, $l_{0}=0.0,0.5$, and $m=0.1,1.0$. The conclusion from this figure is that the thermalization time increases with temperature. This is a feature of the Schwinger model that differs from the expectation of quarkonia in QGP PhysRevD.110.074040 . For low temperatures we have a faster increase of $\mathcal{T}$ with increasing $T$. Above a certain temperature, we enter a linear regime. Here the second term of Eq. (12) is negligible compared to the first term which is independent of $T$. The figure also reassures the previous results in that at any fixed $T$ for a wide range of $T$, increasing $D,l_{0}$ or $m$ results in a larger $\mathcal{T}$.
@@
Figure 7: Thermalization time $\mathcal{T}$ as a function of the environment’s temperature $T$ for $N=12$, $D=2,5$, $l_{0}=0.0,0.5$, $m=0.1,1.0$. The error bars, which are not visible due to the y-scale of the plot, are estimated to be $0.1$ as inferred from the data in table 1 and discussed in section 4.4. The lines are drawn solely for guidance.
@@
Additionally, in figure 9 we present results for the subtracted mutual information $\Delta I$ as a function of time at $D=5.0$, $m=1.0$, $l_{0}=0.5$, for various temperature values over the range $T\in[7,100]$. The figure shows an increasing peak value for $\Delta I$ with increasing $T$. Further, at late times, an increasing $T$ makes the $\Delta I$ tend to the steady state value faster. Both effects seem to be subtle but might be attributed to the following. Higher peaks at higher $T$ suggests that increasing $T$ allows the string’s charges to build more mutual information in the transient time regime. The subtracted kinetic energy $\Delta K$ comparison between $T=7$ and $T=100$ shown in figure 8$(a)$ provides numerical evidence that higher $T$ gives a smaller $\Delta K$ peak, which in turn implies the charges might be spatially closer for a longer time as $T$ increases. Increasing $T$ also increases the particle number production that eventually can weaken the string tension through the creation of screening charges and destroy the mutual information faster. This would explain the late time pattern. Numerical evidence for the particle production levels are presented in figure 8$(b)$, where we plot the particle number for the case when the initial state is the Dirac vacuum with the string present to compare between $T=7$ and $T=100$. We call this observable $\mathcal{P}$ and define it mathematically as
@@
The subfigures $(c)$, $(d)$ in figure 8 show the subtracted electric field $\Delta F(n)$ as a function of the link number $n$ and time $t$ for $T=7,100$ respectively. From these we can qualitatively see that around $t=50$ the links 3 and 8 for $T=100$ have a larger $\Delta F(n)$ compared to $T=7$ which implies the string expands more with larger $T$. This can be related to figure 8$(a)$ where the $T=100$ case has a slower late time decrease to the steady state value of $\Delta K$ and this crossover occurs around $t=50$, when $\Delta K$ for $T=7$ goes below that of $T=100$. Figure 8$(d)$ also shows the generally longer thermalization time that occurs with higher temperature as the middle links have a higher $\Delta F(n)$ for late times compared to figure 8$(c)$.
@@ 5 Conclusion and outlook
The electric field observable on the middle lattice link was used to define a thermalization time $\mathcal{T}$. A similar definition using the total energy of the system in the case of the Schwinger boson has shown consistent pattern of results for the dependence of the thermalization time on $D$. Using the electric field definition, we have observed for both the string and the Schwinger boson initial states that $\mathcal{T}$ increases with increasing dissipator strength $D$, which determines the interaction strength between the system and the environment. Further, for the case of the string, $\mathcal{T}$ has also been observed to increase with increasing applied background electric field $l_{0}$, fermion mass $m$, and environment temperature $T$.
@@ Appendix B Adaptive Time-Dependent DMRG
The Lindblad operator $\mathcal{L}$ of Eq. (16) is partitioned in three groups which we call even, odd and Taylor as shown in figure 16. The even group (bottom green boxes) includes four-site operators that start at sites 0, 4 etc. The odd group (top blue box) includes four-site operators that start at sites 2, 6 etc. Any other term from $\mathcal{L}$ that does not fit into these two groups is put into the Taylor group. The exponentiation of this group as required in time evolution shown in Eq. (20) is done using the Taylor expansion approximation. We can then express the Taylor expansion operator as an MPO (middle red box).
@@
Figure 16: Schematic drawing example of the even group (bottom green boxes), Taylor group (middle red box) and odd group (top blue box) for $\mathcal{L}$ in Eq. (16) at $N=4$. These groups are applied to the MPS at the very bottom representing $\rho_{S}$ as required by Eq. (20), although here only $e^{\tau\mathcal{L}_{O}}e^{\frac{\tau}{2}\mathcal{L}_{T}}e^{\frac{\tau}{2}\mathcal{L}_{E}}$ is shown.

### v2

@@ 1 Introduction
Despite these advances, several open questions remain regarding the impact of gauge DOF truncation, the role of external fields in string dynamics, and the dynamics of the theory’s stable particle, called the Schwinger boson. Prior studies often truncate the gauge DOF Hilbert space to allow only a small number of states per link, potentially limiting the proximity to the continuum limit. Additionally, the influence of an applied background electric field on string behavior in open quantum systems (OQS) has not been explored.
@@ 2 Theoretical framework of the open lattice Schwinger model
An open quantum system is described by the system of interest $S$ and its environment $E$ through the Hamiltonians $H_{S}$, $H_{E}$ and $H_{I}$, where the latter is the interaction Hamiltonian between $S$ and $E$ Breuer:2002pc ; Rivas:2012ugu . It is usually the case that we are interested in measuring observables only on $S$ but nevertheless want to study the effect of $E$ on $S$. Hence, we only aim to follow the dynamics of the reduced density matrix for the system $\rho_{S}$ defined by tracing out the degrees of freedom of $E$.
@@
where $\rho_{E}=e^{-\beta H_{E}}/\text{Tr}(e^{-\beta H_{E}})$ is the environment’s density matrix in the Gibbs state. Further, we work in the quantum Brownian motion (QBM) limit given by the separation of time scales $\tau_{R}\gg\tau_{E}$, $\tau_{S}\gg\tau_{E}$. The first inequality implies that the system with relaxation time $\tau_{R}\sim T/H_{I}^{(\text{int})^{2}}$ relaxes much more slowly than the environment’s correlation time $\tau_{E}\sim 1/T$, where $H_{I}^{(\text{int})}$ is the interaction Hamiltonian in the interaction picture. We can interpret $H_{I}^{(\text{int})}$ in the definition of $\tau_{R}$ as the gap between the ground and first excited state of $H_{I}^{(\text{int})}$. The second inequality suggests that the intrinsic time scale of the system $\tau_{S}\sim 1/H_{S}$ is much greater than $\tau_{E}$, which is valid in the limit $T\gg H_{S}$. Here we can interpret $H_{S}$ to be the spectral gap of $H_{S}$ which is $1/\sqrt{\pi}$ at $m=0$ in the continuum limit adam . These limits further justify the assumption that $\rho_{E}$ remains for all times in the thermal Gibbs state defined by $H_{E}$ as any perturbation on the environment’s Gibbs state will decay too fast to affect the system which we are interested in. The high temperature limit we are using, where the environment is always in thermal equilibrium, is relevant for the case of a hot nuclear QGP medium oqs_for_quarkonia . Under these assumptions, the evolution of the density matrix of $S$ is given by the Lindblad master equation
@@ 3 Lindbladian evolution scheme with tensor networks
We represent the right-hand side of Eq. (19) as a global MPO with two singular value cutoffs $\epsilon_{1}$, $\epsilon_{2}$. Using the ITensors Julia library ITensors allows to convert $\mathcal{L}_{T}$ automatically to an MPO and can apply a truncation on the singular values in the process, which we call $\epsilon_{1}$. The ratio of the sum of squares of the neglected singular values to the sum of squares of all singular values will not exceed this cutoff at any given singular value decomposition truncation. Then $\epsilon_{2}$ is used when we are multiplying $\mathcal{L}_{T}$ onto itself to form powers of this operator as needed for the right hand side of Eq. (19). Finally, we use the adaptive time-dependent DMRG algorithm Xiang_2023 to apply the above operators with a singular value cutoff $\epsilon$, and evolve our state $\rho_{S}(t)$. For more details on this algorithm and the partition of $\mathcal{L}$ into groups see appendix B. Throughout this paper we fix $D(n-k)=D\delta_{n,k}$ open_schwinger_original , $x=1$, $\epsilon=10^{-11}$, $\epsilon_{1}=\epsilon_{2}=10^{-9}$, $\kappa=2$ and $\tau=0.01$, except from section 4.4 where we set $x=4$, $\tau=0.001$ due to the larger $N$ used. For the high-temperature limit in which we are working, the environment’s correlation length is well approximated by $1/T$, implying that it is small. Hence, the delta function dissipator provides a good description for the environment in this regime.
The above methods allow us to simulate the time evolution of meson type states according to Eq. (14) and gather results for their dynamics in a hot medium which we turn to in the next section.
@@ 4.3 Temperature dependence of the string’s thermalization
In figure 7 we plot $\mathcal{T}$ as a function of $T$ for $D=2,5$, $l_{0}=0.0,0.5$, and $m=0.1,1.0$. The conclusion from this figure is that the thermalization time increases with temperature. This is a feature of the Schwinger model that differs from the expectation of quarkonia in QGP PhysRevD.110.074040 . For low temperatures we have a faster increase of $\mathcal{T}$ with increasing $T$. Above a certain temperature, we enter a linear regime. The figure also reassures the previous results in that at any fixed $T$ for a wide range of $T$, increasing $D,l_{0}$ or $m$ results in a larger $\mathcal{T}$.
@@
Figure 7: Thermalization time $\mathcal{T}$ as a function of the environment’s temperature $T$ for $N=12$, $D=2,5$, $l_{0}=0.0,0.5$, $m=0.1,1.0$. The error bars, which are not visible due to the y-scale of the plot, are estimated to be $0.1$ as inferred from the data in table 1 and discussed in section 4.4. The lines represent the fit described by Eq. (23).
In Eq. (A3) of open_schwinger , the authors provide an approximation for the relaxation rate, which exhibits a quadratic dependence on the Lindblad jump operators in the numerator. Expanding this quadratic dependence results in a sum of terms proportional to $1/T^{0}$, $1/T$, and $1/T^{2}$. Consequently, the thermalization time is inversely proportional to this sum, directly explaining the observed increase in thermalization time with temperature in figure 7, as well as the initial nonlinear behavior and the transition to a linear regime. To quantitatively capture this dependence, we fit the data in figure 7 using the function
$f(T)=\frac{a}{(b+c/T)^{2}},$ (23)
with fitting parameters $a,b,c$. The fitted function accurately reproduces the observed functional behavior.
In the case of QCD, the key difference arises from the temperature dependence of $D(k)$. Specifically, $D(k)$ scales as $T^{3}$, which leads to an increase in the relaxation rate with temperature. This behavior is due to the relationship between $D(k)$ and the heavy quark diffusion coefficient $\kappa$ oqs_for_quarkonia ; PhysRevD.97.074009, which has been shown to exhibit a $T^{3}$ dependence PhysRevLett.100.052301.
@@
In figure 9 we present results for the subtracted mutual information $\Delta I$ as a function of time at $D=5.0$, $m=1.0$, $l_{0}=0.5$, for various temperature values over the range $T\in[7,100]$. The figure shows an increasing peak value for $\Delta I$ with increasing $T$. Further, at late times, an increasing $T$ makes the $\Delta I$ tend to the steady state value faster. Both effects seem to be subtle but might be attributed to the following. Higher peaks at higher $T$ suggests that increasing $T$ allows the string’s charges to build more mutual information in the transient time regime. The subtracted kinetic energy $\Delta K$ comparison between $T=7$ and $T=100$ shown in figure 8$(a)$ provides numerical evidence that higher $T$ gives a smaller $\Delta K$ peak, which in turn implies the charges might be spatially closer for a longer time as $T$ increases. Increasing $T$ also increases the particle number production that eventually can weaken the string tension through the creation of screening charges and destroy the mutual information faster. This would explain the late time pattern. Numerical evidence for the particle production levels are presented in figure 8$(b)$, where we plot the particle number for the case when the initial state is the Dirac vacuum with the string present to compare between $T=7$ and $T=100$. We call this observable $\mathcal{P}$ and define it mathematically as
@@
The subfigures $(c)$, $(d)$ in figure 8 show the subtracted electric field $\Delta F(n)$ as a function of the link number $n$ and time $t$ for $T=7,100$ respectively. From these we can qualitatively see that around $t=50$ the links 3 and 8 for $T=100$ have a larger $\Delta F(n)$ compared to $T=7$ which implies the string expands more with larger $T$ at later times. This can be related to figure 8$(a)$ where the $T=100$ case has a slower late time decrease to the steady state value of $\Delta K$ and this crossover occurs around $t=50$, when $\Delta K$ for $T=7$ goes below that of $T=100$. Figure 8$(d)$ also shows the generally longer thermalization time that occurs with higher temperature as the middle links have a higher $\Delta F(n)$ for late times compared to figure 8$(c)$.
@@ 5 Conclusion and outlook
The electric field observable on the middle lattice link was used to define a thermalization time $\mathcal{T}$. A similar definition using the total energy of the system in the case of the Schwinger boson has shown a consistent pattern of results for the dependence of the thermalization time on $D$. Using the electric field definition, we have observed for both the string and the Schwinger boson initial states that $\mathcal{T}$ increases with increasing dissipator strength $D$, which determines the interaction strength between the system and the environment. Further, for the case of the string, $\mathcal{T}$ has also been observed to increase with increasing applied background electric field $l_{0}$, fermion mass $m$, and environment temperature $T$.
@@ Appendix B Adaptive Time-Dependent DMRG
The Lindblad operator $\mathcal{L}$ of Eq. (16) is partitioned in three groups which we call even, odd and Taylor as shown in figure 16. The even group (bottom green boxes) includes four-site operators that start at sites 0, 4 etc. The odd group (top blue box) includes four-site operators that start at sites 2, 6 etc. Any other term from $\mathcal{L}$ that does not fit into these two groups is put into the Taylor group. The exponentiation of this group as required in time evolution shown in Eq. (20) is done using the Taylor expansion approximation. We can then express the Taylor expansion operator as an MPO (middle red boxes).
@@
Figure 16: Schematic drawing example of the even group (bottom green boxes), Taylor group (middle red boxes) and odd group (top blue box) for $\mathcal{L}$ in Eq. (16) at $N=4$. These groups are applied to the MPS at the very bottom representing $\rho_{S}$ as required by Eq. (20), although here only $e^{\tau\mathcal{L}_{O}}e^{\frac{\tau}{2}\mathcal{L}_{T}}e^{\frac{\tau}{2}\mathcal{L}_{E}}$ is shown.